# Post-detection inference for sequential changepoint localization

### v1

@@ 1 $r_{1}=r_{2}=\cdots=r_{\tau}=1$
4             Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}^{*}}$, define $\tau_{i}$ as the time when $\mathcal{A}$ detects a change, or $\tau$, whichever occurs first
@@
7      Define $r_{t}=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$ for $t=1,\dots,\tau$;
@@ 24            $\mathcal{C}=\mathcal{C}\cup\{t\}$
If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, $\mathcal{C}_{2}$ satisfies the following guarantee for any $\theta_{i}\in\Theta_{i}(i=0,1)$:
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\gamma-\delta,\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta},\text{ when }\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
@@ 4.2 Computationally and statistically efficient subroutines
However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$ and $t_{2}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$. Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient. E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
$U_{t,L}^{j}:=\begin{cases}-\infty&\text{if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),&\text{if}t\leq t_{2}^{j}\leq L\\ \infty&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (4.10)
@@
$U_{t,\infty}^{j}:=\begin{cases}\displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),&\text{if }t\leq t_{2}^{j}\text{ and }t_{1}^{j}<\infty\\ -\infty&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty.\end{cases}$ (4.11)
@@
$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (4.12)
@@ 4.3 A concrete example: Gaussian mean shift problem
We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$.
In this setting, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$
@@
We fix $t\in\{2,\cdots,\tau\}$. Note that we can consider $\theta^{\prime\prime}=0$ and $G(x,\theta)=x+\theta$. Therefore, only need to generate the sequence $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$. For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, we define
@@
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms as we illustrate in Remark 4.5.
For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
$U_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
@@
$-2\log L_{i}\left(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\right)=$
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{ constant },\text{ if }i<t,\\ \text{ constant},\text{ if }i=t,\\ \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{ constant },\text{ if }i>t,\end{cases}$ (4.16)
over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (4.3) is a quadratic function of $\theta_{1}-\theta_{0}$. Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
@@
$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\}$ (4.17)
[…]
@@
$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
Similarly, it is easy to check that for $\mathcal{P}_{0}=\{N(\theta,1):\theta\geq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\leq b\}$, for some $b\leq a$,
$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
@@
$\mathbb{P}_{\theta_{0},\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)<A\right]$
@@
The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology. It is worth noting that $\mathcal{C}^{\prime\prime}$ in (4.12) is computable for parametric families if the following conditions are met:
@@
    The coupling function $G$ is available (which is always the case when distributions are continuous).
@@
    We can construct a confidence interval (and hence confidence sequence) for the parameter of interest.
@@
For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions.

### v2

@@ 1 $r_{1}=r_{2}=\cdots=r_{\tau}=1$
4             Draw $\{X_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta_{0}^{*}}$, define $\tau_{i}$ as the time when $\mathcal{A}$ detects a change, or $\tau$, whichever is first
@@
7      Define $r_{t}=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$ for $t=1,\dots,\tau$;
@@ 24            $\mathcal{C}=\mathcal{C}\cup\{t\}$
If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, we have that for any $\theta_{i}\in\Theta_{i}(i=0,1)$:
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\gamma-\delta,\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta+\gamma}{1-\delta}$
[…]
@@ 4.2 Computationally and statistically efficient subroutines
Note that, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta,\theta^{\prime})\}_{n}$ across all $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Section B.2 for the Poisson distribution.
@@
However, it is often straightforward to compute $t_{1}^{j}=\displaystyle\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$ and $t_{2}^{j}=\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}$. Also, we can often derive a suitable upper bound on the threshold, which does not require any approximation and is computationally and statistically efficient. E.g., we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$. Again, we bound $\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta,\theta^{\prime})$ by $U_{t,L}^{j}$, where for $L<\infty$,
$U_{t,L}^{j}:=\begin{cases}-\infty,\leavevmode\nobreak\ \leavevmode\nobreak\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (4.10)
@@
$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (4.11)
@@
$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (4.12)
@@
In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with Poisson distribution in Appendix B.
@@ 4.3 A concrete example: Gaussian mean shift problem
We consider the most common changepoint model, where pre-change and post-change models are $\mathcal{P}_{i}=\{N(\theta,1):\theta\in\Theta_{i}\}$, for $i=0,1$. Then, we have the following $1-\gamma$ confidence interval for $\theta_{0}$ for each $t\geq 2$:
@@
We fix $t\in\{2,\cdots,\tau\}$. Note that we can consider $\theta^{\prime\prime}=0$ and $G(x,\theta)=x+\theta$. Therefore, only need to generate $\{\epsilon_{n}^{j}\}_{n\in\mathbb{N}}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ for $j=1,\cdots,B$. For $\theta\in\mathcal{S}_{t-1},\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}$, define
@@
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark 4.5. For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
$\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\displaystyle\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}(\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}{L_{t}(\bar{\epsilon}_{1:t-1}^{j}+\theta_{0},\bar{\epsilon}_{t:t^{\prime}}^{j}+\theta_{1};Y_{1}^{j}(\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime}))}$
@@
$-2\log L_{i}\Big{(}\bar{Y}_{1:i-1}^{j}(\theta,\theta^{\prime}),\bar{Y}_{i:t^{\prime}}^{j}(\theta,\theta^{\prime});Y_{1}^{j}(\theta,\theta^{\prime}),\dots,Y_{t^{\prime}}^{j}(\theta,\theta^{\prime})\Big{)}=$
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (4.16)
over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Here, constant means that it does not involve $\theta,\theta^{\prime}$. Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (4.3) is a quadratic function of $\theta_{1}-\theta_{0}$. Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
@@
$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (4.17)
$\tau_{\text{wsr-ripr}}=\inf\left\{\tau\in\mathbb{N}:\sum_{j=1}^{\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (4.18)
@@
$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
Similarly, one can check that for $\mathcal{P}_{0}=\{N(\theta,1):\theta\geq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\leq b\}$, for $b\leq a$,
$t_{1}^{j}=\inf_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\inf\mathcal{S}_{t-1},\inf\mathcal{S}^{\prime}_{\tau-t+1}}^{j}\text{ and }$
$t_{2}^{j}=\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}\tau_{\theta,\theta^{\prime}}^{j}=\tau_{\sup\mathcal{S}_{t-1},\sup\mathcal{S}^{\prime}_{\tau-t+1}}^{j}.$
@@
$\mathbb{P}_{\theta_{0},\infty}[\tau_{\text{wcs-ripr}}\geq t]$
$=\mathbb{P}_{\theta_{0},\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)<A\right]$
@@
The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology. Note that $\mathcal{C}^{\prime\prime}$ in (4.12) is computable if the following conditions are met:
@@
    The coupling function $G$ is available (which is always the case when distributions are continuous). For discrete distributions, even if a suitable $G$ is not available, our method may still be implementable, as demonstrated in Section B.2 for the Poisson case).
@@
    We can construct a confidence interval (and hence confidence sequence) for $\theta$.
@@
For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions. Another concrete example with the Poisson distribution is in Appendix B.